OpenStudy (anonymous):

i really don't understand the process of how to factor this (3x^2+11)^2-16x^2

OpenStudy (anonymous):

\[\diamondsuit^2-\spadesuit^2=(\diamondsuit+\spadesuit)(\diamondsuit -\spadesuit)\]

OpenStudy (anonymous):

what do those symbols mean?

geerky42 (geerky42):

\(a^2-b^2 = (a+b)(a-b)\) In your problem, \(a=3x^2+11\) and \(b = 4x\)

OpenStudy (anonymous):

replace \(\diamondsuit\) by \(3x^2+11\) and \(\spadesuit \) by \(4x\)

OpenStudy (kl0723):

is just a representation of a variable just like @geerky42 explained same thing :0

geerky42 (geerky42):

Do these explanations help? @torituller97

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

so how do i set that up to solve

OpenStudy (kainui):

The most important part of the entire process is that you use your mind. Think about it, try it out, and figure it out. The only person that can make it make sense to you is you, yourself. @torituller97

OpenStudy (kainui):

One possible way to help yourself understand it is by creating your own simpler problem and solving it. But there are many things you can do to reach understanding.

OpenStudy (anonymous):

but i just don't know how to make sense of it @Kainui

OpenStudy (loser66):

\[\color{red}{a}^2-\color{blue}{b}^2 = (\color{red}{a}+\color{blue}{b})(\color{red}{a}-\color{blue}{b})\] \[(\color{red}{3x^2+11})^2-(\color{blue}{16x^2})^2 = ??\] Imitate the formula

OpenStudy (kainui):

Here write down this formula, go here and this will help you think about it more: http://tinyurl.com/probsolvin

OpenStudy (anonymous):

that link didn't work... but is the formula (a+b)(a^2+ab+b^2)?

geerky42 (geerky42):

That's for sum of cubes.

OpenStudy (anonymous):

@Kainui

geerky42 (geerky42):

\(a^3+b^3 =(a+b)(a^2+ab+b^2) \)

OpenStudy (anonymous):

what am i trying to find? @geerky42

OpenStudy (anonymous):

actually \[a^3+b^3 =(a+b)(a^2\color{red}-ab+b^2)\]

geerky42 (geerky42):

Ya, but it's irrelevant anyway. We just need formula for difference of squares.

OpenStudy (anonymous):

can you show me how to set this up because i don't know what a and b are?